OpenStudy (anonymous):

A lab has 65 mice. Some are white and the rest are gray. Which of the following could be the ratio of white mice to gray mice in the lab? 1 : 3 5 : 8 5 : 11 3 : 4 5 : 12

OpenStudy (anonymous):

0.o none? I think. none add up to 65.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

1:3

OpenStudy (diyadiya):

5:8 5/8=25/40 25+40=65

OpenStudy (anonymous):

because 1/3 x 15/15 = 15/45 and so 15+45= 60

OpenStudy (anonymous):

gracias

OpenStudy (anonymous):

Got it ????

OpenStudy (diyadiya):

it isn't 60 its 65

OpenStudy (anonymous):

5:12

OpenStudy (anonymous):

If x is a positive integer then 2(x)3 =